# OODFace: Benchmarking Robustness of Face Recognition under Common Corruptions and Appearance Variations

### v1

@@ B.1 LFW-C
For LFW-C, we design 5 levels for each corruption, ranging from level 1 (mild) to level 5 (extreme). In this section, we provide the specific data tables corresponding to levels 1-5, as shown in Tab. H.1, Tab. H.2, Tab. H.3, Tab. H.4, Tab. H.5. For each model, we present a radar chart that shows the performance of the model across 20 different types of corruption, providing a clearer view of how the model’s performance varies under different types of corruption in each category, as shown in Fig. H.1. For each corruption category, we also provide line charts showing the model’s performance across the 5 levels of each corruption, illustrating the performance fluctuations as the level of corruption increases, as shown in Fig. H.3.
@@ B.2 LFW-V
Similarly, for LFW-V, we design 5 levels for each variation. We supplement the specific data tables corresponding to levels 1-5 in this section, as shown in Tab. H.6, Tab. H.7, Tab. H.8, Tab. H.9, Tab. H.10. We present the radar charts in Fig. H.2, as well as the line charts showing the 5 levels for each category, in Fig. H.4.
@@ Appendix C Results on CFP-C/V
For CFP, we follow the experimental setup used for LFW-C/V and test 19 FR models on 20 types of common corruptions and 10 types of appearance variations, categorized into Open-source Model Eval, Architecture Eval, and Loss Function Eval. The test results are shown in Tab. H.11 and Tab. H.17. We also report the Relative Corruption Error (RCE), as shown in Fig. H.5, for each model across different corruption categories. Additionally, in Appearance Variations, we further illustrate the Relative Variation Error (RVE), as shown in Fig. H.6, for each model across variation categories.
We also provide the specific data tables corresponding to levels 1-5, as shown in Tab. H.12, Tab. H.13, Tab. H.14, Tab. H.15, Tab. H.16 and Tab. H.18, Tab. H.19, Tab. H.20, Tab. H.21, Tab. H.22. We present the radar charts and line charts showing the 5 levels for each category, in Fig. H.7, Fig. H.8 and Fig. H.9, Fig. H.10.
@@ Appendix D Results on YTF-C/V
For YTF, we also provide the average accuracy data on YTF-C and YTF-V, shown in Tab. H.23 and Tab. H.29, respectively. We similarly report the Relative Corruption Error (RCE), as shown in Fig. H.11, and the Relative Variation Error (RVE), as shown in Fig. H.12. Additionally, we provide the specific data tables corresponding to levels 1-5, as shown in Tab. H.24, Tab. H.25, Tab. H.26, Tab. H.27, Tab. H.28 and Tab. H.30, Tab. H.31, Tab. H.32, Tab. H.33, Tab. H.34. Furthermore, we present the radar charts and line charts showing the 5 levels for each category, in Fig. H.13, Fig. H.14 and Fig. H.15, Fig. H.16.
For the YTF data, our findings are as follows: First, the FR models’ clean accuracy on YTF is lower than on LFW and CFP. This is because YTF contains more complex data from various scenes, which is reflected in the final results. As shown in Tab. H.23, unlike LFW-C and CFP-C, the differences between models on YTF-C are more pronounced. For example, ArcFace achieves the highest clean accuracy and robustness in the Data & Processing category, while TransFace demonstrates advantages in specific categories like Snow, Color Shift, and Motion Blur.
On YTF-V, due to YTF’s complex scenes, the performance drop caused by appearance variations is more severe, especially for ArcFace and ElasticFace. Although they achieve higher clean accuracy, their performance suffers more under variations. On the other hand, AdaFace, despite not having the highest clean accuracy, maintains the best robust accuracy under appearance variations. This finding slightly differs from the conclusions drawn on LFW and CFP.
Figure D.1: Additional results on testing GPT-4 mini for FR.
@@
Appendix E Additional Results on Face Masks
@@
In the field of face recognition (FR), masks typically refer to physical disguises used to deceive or conceal the identity of the wearer, preventing accurate recognition by the system. As an extension, we conduct physical experiments with masks.
@@
In our experiments, we create five different types of masks using various collection methods and fabrication processes. We categorize them from A to E based on their realism, with Mask A being directly collected through photography and created using 2D printing technology with paper material. Masks B and Masks C are created by scanning facial data using 3D scanning technology, with sandstone and paper pulp materials, respectively. Masks D and Masks E are also created using 3D scanning and silicone material, with Masks D and Masks E having higher facial fit and elasticity.
@@
Subsequently, we test these masks on the Common Corruptions and Appearance Variations we design, in order to examine the characteristics of different masks and their impact in OOD scenarios. The results are shown in Tab. H.35, Tab. H.36, Tab. H.37, Tab. H.38, Tab. H.39 and Tab. H.40, Tab. H.41, Tab. H.42, Tab. H.43, Tab. H.44,. The data reveals differences from real faces and exhibits performance degradation trends that differ from those of real faces.
Appendix F Data Augmentation as Potential Defense
We further explore potential defense measures to enhance robustness. Based on the methods described in [52], we employ existing defense strategies available, which can be categorized into two main approaches: input transformation (e.g., R&P [50], Bit-Red [51], JPEG [13]) and adversarial training (e.g., PGD-AT  [32], TRADES [53]). We test 10 robust models based on these approaches, as shown in Tab .A.1. However, as we observe, for input transformation methods, there is no significant impact on the model’s performance. Specifically, Bit-Red [51] increases the average accuracy of Softmax-IR by 0.02, while Bit-Red [51] and JPEG [13] result in slight performance degradation. For PGD-AT [32] and TRADES [53], which are designed for adversarial training, no gains are observed under natural corruptions during testing. On the contrary, due to a notable drop in clean data performance after adversarial training, the overall model performance declines. This indicates that existing defense strategies fail to address Out-of-Distribution (OOD) scenarios, and enhancing the robustness of FR remains an open challenge for future research.
Appendix G VLMs as Potential solutions
In Sec.5.4 of the main text, we mentioned that, to explore better solutions for OOD scenarios in face recognition (FR), we investigated the use of Vision-Language Models (VLMs), a previously unexplored approach. The GPT-4o-mini we tested demonstrated strong FR capabilities, even being able to recognize specific types of corruptions present in images. We provide more examples, as shown in Fig D.1. When testing 50 pairs of faces with added common corruptions on the web version of GPT-4o-mini, all 50 pairs were successfully recognized, indicating the potential of VLMs in addressing the OOD data challenge.
In fact, we attempted to use the GPT API for batch testing of face data, such as LFW. Unfortunately, the API version of GPT-4o-mini refuse to respond to face-related queries, stating that it “Cannot perform face recognition tasks." This may be due to privacy concerns related to tasks involving FR, which are considered sensitive and subject to security alignment policies, leading to the refusal. In our experiments, we also test models such as Claude 3 Sonnet and Gemini, but they do not seem to return correct FR results. Therefore, while the results with GPT-4o-mini suggest that leveraging the generalization capabilities of large models may be a promising solution for handling OOD scenarios, the specific implementation details and usage methods still require further exploration.
@@ Appendix H More Visualization Results
Figure H.1: Graded radar charts for each model on LFW-C.
Figure H.2: Graded radar charts for each model on LFW-V.
Figure H.3: Graded line charts for each corruption on LFW-C.
Figure H.4: Graded line charts for each corruption on LFW-V.
Figure H.5: RCE results on CFP-C.
Figure H.6: RVE results on CFP-V.
Figure H.7: Graded radar charts for each model on CFP-C.
Figure H.8: Graded radar charts for each model on CFP-V.
Figure H.9: Graded line charts for each corruption on CFP-C.
Figure H.10: Graded line charts for each corruption on CFP-V.
Figure H.11: RCE results on YTF-C.
Figure H.12: RVE results on YTF-V.
Figure H.13: Graded radar charts for each model on YTF-C.
Figure H.14: Graded radar charts for each model on YTF-V.
Figure H.15: Graded line charts for each corruption on YTF-C.
Figure H.16: Graded line charts for each corruption on YTF-V.
@@
Table H.1: Accuracy of 19 FR models on LFW-C level 1.
@@
Table H.2: Accuracy of 19 FR models on LFW-C level 2.
@@
Table H.3: Accuracy of 19 FR models on LFW-C level 3.
@@
Table H.4: Accuracy of 19 FR models on LFW-C level 4.
@@
Table H.5: Accuracy of 19 FR models on LFW-C level 5.
@@
Table H.6: Accuracy of 19 FR models on LFW-V level 1.
@@
Table H.7: Accuracy of 19 FR models on LFW-V level 2.
@@
Table H.8: Accuracy of 19 FR models on LFW-V level 3.
@@
Table H.9: Accuracy of 19 FR models on LFW-V level 4.
@@
Table H.10: Accuracy of 19 FR models on LFW-V level 5.
@@
Table H.11: Accuracy of 19 FR models on CFP-C.
@@
Table H.12: Accuracy of 19 FR models on CFP-C level 1.
@@
Table H.13: Accuracy of 19 FR models on CFP-C level 2.
@@
Table H.14: Accuracy of 19 FR models on CFP-C level 3.
@@
Table H.15: Accuracy of 19 FR models on CFP-C level 4.
@@
Table H.16: Accuracy of 19 FR models on CFP-C level 5.
@@
Table H.17: Accuracy of 19 FR models on CFP-V.
@@
Table H.18: Accuracy of 19 FR models on CFP-V level 1.
@@
Table H.19: Accuracy of 19 FR models on CFP-V level 2.
@@
Table H.20: Accuracy of 19 FR models on CFP-V level 3.
@@
Table H.21: Accuracy of 19 FR models on CFP-V level 4.
@@
Table H.22: Accuracy of 19 FR models on CFP-V level 5.
@@
Table H.23: Accuracy of 19 FR models on YTF-C.
@@
Table H.24: Accuracy of 19 FR models on YTF-C level 1.
@@
Table H.25: Accuracy of 19 FR models on YTF-C level 2.
@@
Table H.26: Accuracy of 19 FR models on YTF-C level 3.
@@
Table H.27: Accuracy of 19 FR models on YTF-C level 4.
@@
Table H.28: Accuracy of 19 FR models on YTF-C level 5.
@@
Table H.29: Accuracy of 19 FR models on YTF-V.
@@
Table H.30: Accuracy of 19 FR models on YTF-V level1.
@@
Table H.31: Accuracy of 19 FR models on YTF-V level2.
@@
Table H.32: Accuracy of 19 FR models on YTF-V level3.
@@
Table H.33: Accuracy of 19 FR models on YTF-V level4.
@@
Table H.34: Accuracy of 19 FR models on YTF-V level5.
@@
Table H.35: Accuracy of 8 open-source models on Mask-A of corruptions.
@@
Table H.36: Accuracy of 8 open-source models on Mask-B of corruptions.
@@
Table H.37: Accuracy of 8 open-source models on Mask-C of corruptions.
@@
Table H.38: Accuracy of 8 open-source models on Mask-D of corruptions.
@@
Table H.39: Accuracy of 8 open-source models on Mask-E of corruptions.
@@
Table H.40: Accuracy of 8 open-source models on Mask-A of variations.
@@
Table H.41: Accuracy of 8 open-source models on Mask-B of variations.
@@
Table H.42: Accuracy of 8 open-source models on Mask-C of variations.
@@
Table H.43: Accuracy of 8 open-source models on Mask-D of variations.
@@
Table H.44: Accuracy of 8 open-source models on Mask-E of variations.

### v2

@@ A.2 Implementation Details of Variations
Common Corruptions Appearance Variations FR Models Extensions
Sensor Movement Occlusion L & W D & P Age Rotation Facial Exp. Accessories Categories Num. Open-source API Defense VLMs Phys. Exp.
[…]
Table A.3: Comparison with related work. Our benchmark provides the most comprehensive evaluation.
@@
Variation Levels: For age adjustments, the levels indicate the degree of age reduction or increase. For facial expressions, the levels represent the extent of mouth and eye movements. Head pose levels denote the degree of facial rotation. For accessory additions, the levels are defined as follows: glasses (clear lenses), bang occlusion, glasses (reflective lenses), glasses (clear lenses) + bang, and glasses (reflective lenses) + bang. For the makeup variation, we adopt the styles defined in [27]: Japanese style, Korean style, Retro style, Flashy style, and Smoky-eyes style. Each level represents an increase in perceived appearance interference.
A.3 Comparison with Related Work
As mentioned in Section 2.2, related studies [2, 12, 3, 39] have also explored data corruption in face recognition. Compared to them, our benchmark is more comprehensive in terms of OOD types, evaluation datasets, and the FR models studied, as shown in Tab. A.3.
For instance, [3] examines 5 Lighting & Weather OOD categories and evaluates 3 open-source models; [39] primarily focuses on 9 OOD challenges under the Occlusion category, testing 8 open-source models; [2] centers on variations in Facial Expression, while [12] covers a broader range of common corruptions, including Sensor, Lighting & Weather, and Data & Processing, alongside discussions on defense strategies. However, our benchmark systematically evaluates the most extensive set of OOD scenarios, categorized into Common Corruptions and Appearance Variations, covering nine main categories and 30 subcategories. We assess 19 open-source models, 3 commercial APIs, and further investigate defense mechanisms, physical-world face mask experiments, and the potential utilization of VLMs.
Notably, previous works have not comprehensively considered corruption scenarios specific to FR. In contrast, we are the first to investigate such OOD cases within a unified robustness benchmark, including OOD challenges like Facial Distortion, Random Occlusion, Age, and Accessories—factors explicitly designed for FR yet previously unexplored.
A.4 Naturalness of OOD Synthesis.
While it is impossible to exhaust all real-world OOD types, we systematically designed 30 OOD categories into 5 levels, to serve as a practical testbed for controllable robustness evaluation. Certain OOD types originate from digital artifacts, such as noise corruptions and JPEG compression, while others follow the synthesis strategy in [22], where we carefully adjust parameters to ensure realism (details in Appendix A.1). Specifically, for Appearance Variations, we employ state-of-the-art generative simulation methods proven to closely approximate real facial variations [42, 28, 27]. Despite the inevitable gap between synthetic and real data, our experiments demonstrate that model performance under synthetic conditions aligns well with real-world results. Quantitative analysis is provided in Tab. A.4.
Quantitative Analysis. Firstly, we focus on one of the most complex variations—age progression—as an example, by leveraging the AgeDB [38] dataset. We analyze FR model performance under both synthetically generated and real-world age variations. AgeDB provides identity-labeled age variations, allowing us to evaluate the four FR models introduced in Sec. 4.4 under both conditions, results are shown in Tab. A.4. Notably, model performance remains consistent across synthetic and real-world aging scenarios, validating the effectiveness of our proposed OOD benchmark for assessing FR robustness.
Data Quality Verification. Data quality verification is also a critical part of our benchmark. During OOD scenario simulation, we ensure data integrity by carefully adjusting the severity levels of each corruption and variation to maintain human-recognizable faces. Detailed verification procedures are provided in Appendix A.1 and A.2.
ElasticFace AdaFace TransFace TopoFR
Synthetic Age- 96.32 96.48 96.12 96.08
Age+ 96.39 96.6 96.39 96.19
Real Young 95.00 95.50 94.50 94.00
Old 96.50 97.00 96.50 96.00
Table A.4: Comparison of simulation methods and real data on age variations. The results demonstrate that models exhibit similar trends across both data types.
Figure A.2: Display of face masks and robotic arm for data collection
@@ B.1 LFW-C
For LFW-C, we design 5 levels for each corruption, ranging from level 1 (mild) to level 5 (extreme). In this section, we provide the specific data tables corresponding to levels 1-5, as shown in Tab. I.1, Tab. I.2, Tab. I.3, Tab. I.4, Tab. I.5. For each model, we present a radar chart that shows the performance of the model across 20 different types of corruption, providing a clearer view of how the model’s performance varies under different types of corruption in each category, as shown in Fig. I.1. For each corruption category, we also provide line charts showing the model’s performance across the 5 levels of each corruption, illustrating the performance fluctuations as the level of corruption increases, as shown in Fig. I.3.
@@ B.2 LFW-V
Similarly, for LFW-V, we design 5 levels for each variation. We supplement the specific data tables corresponding to levels 1-5 in this section, as shown in Tab. I.6, Tab. I.7, Tab. I.8, Tab. I.9, Tab. I.10. We present the radar charts in Fig. I.2, as well as the line charts showing the 5 levels for each category, in Fig. I.4.
Variations Evaluation for APIs. Additionally, as shown in Tab. B.1, for LFW-V, no detection rejection occurs as seen in LFW-C, so we report the direct accuracy results. Appearance variations also cause performance degradation, but the impact is smaller compared to corruptions. Among the APIs, Tencent achieves the highest average accuracy, while Aliyun performs better in the Age and Facial Expression. The robustness of these services remains highly correlated with their clean performance, aligning with the conclusions in Sec. 5.2.
Models Accuracy
Variations Aliyun iFLYTEK Tencent
None (clean) 99.65 97.99 99.75
Age Age- 96.60 93.47 96.45
Age+ 96.60 93.50 96.59
Facial Expression Mouth-close 96.67 93.80 96.47
Mouth-open 96.55 93.68 96.55
Eye-close 96.71 93.23 96.64
Eye-open 96.67 94.10 96.72
Rotation Rotation-left 96.67 94.23 96.71
Rotation-right 96.71 94.10 96.67
Accessories Bangs&Glasses 98.06 93.91 98.85
Makeup 99.13 96.60 99.28
Average 97.27 94.42 97.33
Table B.1: Accuracy of 3 commercial APIs on LFW-V.
@@ Appendix C Results on CFP-C/V
For CFP, we follow the experimental setup used for LFW-C/V and test 19 FR models on 20 types of common corruptions and 10 types of appearance variations, categorized into Open-source Model Eval, Architecture Eval, and Loss Function Eval. The test results are shown in Tab. I.11 and Tab. I.17. We also report the Relative Corruption Error (RCE), as shown in Fig. I.5, for each model across different corruption categories. Additionally, in Appearance Variations, we further illustrate the Relative Variation Error (RVE), as shown in Fig. I.6, for each model across variation categories.
We also provide the specific data tables corresponding to levels 1-5, as shown in Tab. I.12, Tab. I.13, Tab. I.14, Tab. I.15, Tab. I.16 and Tab. I.18, Tab. I.19, Tab. I.20, Tab. I.21, Tab. I.22. We present the radar charts and line charts showing the 5 levels for each category, in Fig. I.7, Fig. I.8 and Fig. I.9, Fig. I.10.
@@ Appendix D Results on YTF-C/V
For YTF, we also provide the average accuracy data on YTF-C and YTF-V, shown in Tab. I.23 and Tab. I.29, respectively. We similarly report the Relative Corruption Error (RCE), as shown in Fig. I.11, and the Relative Variation Error (RVE), as shown in Fig. I.12. Additionally, we provide the specific data tables corresponding to levels 1-5, as shown in Tab. I.24, Tab. I.25, Tab. I.26, Tab. I.27, Tab. I.28 and Tab. I.30, Tab. I.31, Tab. I.32, Tab. I.33, Tab. I.34. Furthermore, we present the radar charts and line charts showing the 5 levels for each category, in Fig. I.13, Fig. I.14 and Fig. I.15, Fig. I.16.
For the YTF data, our findings are as follows: First, the FR models’ clean accuracy on YTF is lower than on LFW and CFP. This is because YTF contains more complex data from various scenes, which is reflected in the final results. As shown in Tab. I.23, unlike LFW-C and CFP-C, the differences between models on YTF-C are more pronounced. For example, ArcFace achieves the highest clean accuracy and robustness in the Data & Processing category, while TransFace demonstrates advantages in specific categories like Snow, Color Shift, and Motion Blur.
On YTF-V, due to the more complex scenes of YTF, the performance drop caused by variations is more severe, especially for ArcFace and ElasticFace. Although they achieve higher clean accuracy, their performance suffers more under variations. On the other hand, AdaFace, despite not having the highest clean accuracy, maintains the best robust accuracy under appearance variations. This finding slightly differs from the conclusions drawn on LFW and CFP.
Figure D.1: Testing GPT-4o mini for FR. GPT-4o mini demonstrates robust FR capabilities.
Figure D.2: Additional results on testing GPT-4o-mini for FR.
@@
Appendix E Exploration of Potential Defense
We further explore potential defense measures to enhance robustness. Based on the methods described in [56], we employ existing defense strategies available, which can be categorized into two main approaches: input transformation (e.g., R&P [54], Bit-Red [55], JPEG [15]) and adversarial training (e.g., PGD-AT  [35], TRADES [57]). We test 10 robust models based on these approaches, as shown in Tab .A.1. However, as we observe, for input transformation methods, there is no significant impact on the model’s performance. Specifically, Bit-Red [55] increases the average accuracy of Softmax-IR by 0.02, while Bit-Red [55] and JPEG [15] result in slight performance degradation. For PGD-AT [35] and TRADES [57], which are designed for adversarial training, no gains are observed under natural corruptions during testing. On the contrary, due to a notable drop in clean data performance after adversarial training, the overall model performance declines.
Furthermore, we explored several advanced restoration methods [51, 63, 31], based on GANs, Transformers, and Diffusion models, as potential defense strategies against OOD scenarios. Our experiments yielded interesting results, as summarized in Tab. A.2, where we report the changes in OOD robust accuracy for different FR models after applying these defense methods.
From the Perspective of OOD Category: We observe that for certain OOD categories, the FR model’s recognition performance improved after applying restoration methods, particularly for noise-related corruptions under the Data & Processing category. This is likely because restoration methods are primarily designed for denoising and deblurring, leading to notable improvements in handling noise-related corruptions, as shown in the left panel of Fig. 8. However, for other OOD categories such as Lighting & Weather and Occlusion, recognition performance actually degraded. This could be due to restoration methods distorting key facial landmarks or interfering with feature extraction, as illustrated in the right panel of Fig. 8.
From the Perspective of FR Models: An interesting phenomenon we observed is that models achieving the best performance in OOD testing, such as AdaFace, exhibited little to no improvement across all OOD categories after applying restoration-based defenses. This may be due to AdaFace’s inherent robustness, which already enables it to mitigate noise and other perturbations. However, the restoration methods inadvertently distorted some of the original facial landmarks, disrupting AdaFace’s feature extraction process. As a result, the performance degradation caused by these distortions outweighed the potential benefits of restoration, ultimately leading to a decline in FR accuracy. Conversely, for models that are more vulnerable to OOD challenges, such as SphereFace, restoration-based defenses led to improvements in most categories, with a substantial increase in overall accuracy. The Data & Processing category, which poses the greatest threat to FR performance, presented significant challenges for less robust models. In these cases, the removal of noise through restoration methods provided a notable performance gain, effectively enhancing the model’s robustness.
Therefore, employing restoration methods as a defense strategy could be a promising solution for addressing OOD scenarios, particularly for noise-related corruptions that pose significant challenges. However, since existing restoration techniques are primarily designed for denoising and deblurring, their effectiveness remains limited when facing a broader range of real-world OOD challenges, as explored in OODFace. In such cases, the interference caused by restoration methods in the feature extraction process introduces new challenges, leading to performance degradation in originally robust models.
Notably, none of the individual defense methods in our experiments were able to effectively mitigate more than 30% of the OOD categories. This suggests the need for more generalized and precise defense strategies. These results indicate that existing defense strategies are insufficient for effectively handling comprehensive out-of-distribution scenarios. Enhancing the overall robustness of facial recognition models remains an open challenge for future research.
Appendix F VLMs as Potential Solutions
In Sec. 5.4 of the main paper, we explored the potential of Vision-Language Models (VLMs) for addressing FR OOD scenarios, a direction that remains largely unexplored. Our evaluation of GPT-4o-mini revealed strong FR capabilities, including the ability to recognize specific corruption types (see Fig D.2 for examples). Further, we tested both closed-source commercial models (GPT-4o-mini, Qwen-VL-Plus) and open-source models (LLaVA-NeXT-LLaMA3-8B, InternVL2.5-8B) on face images corrupted with 20 common corruptions designed in our benchmark. The prompt used was:
“Determine if the faces in the image belong to the same person. Reply with 1 if they are the same, and 0 otherwise. Your answer must be either 0 or 1."
This setup requires the models to directly output a binary decision on face identity matching. Notably, open-source models fail to make reliable predictions, outputting nearly all 0s, while closed-source models achieve promising results. Qwen-VL-Plus reaches 87.76% accuracy, and GPT-4o-mini achieves an impressive 98.98%.
Although our experimental results suggest that large models’ generalization ability could be a promising solution for handling OOD scenarios, their practical application still faces certain limitations and challenges. Notably, only proprietary models achieve strong performance, raising concerns regarding accessibility and deployment constraints. Additionally, the use of facial data for model training introduces significant security and privacy considerations that must be seriously addressed. Furthermore, understanding why VLMs achieve superior FR performance and how to effectively integrate VLMs into existing FR pipelines to enhance robustness remain important directions for future research.
Appendix G Additional Results on Face Masks
@@
In the field of face recognition, masks typically refer to physical disguises used to deceive or conceal the identity of the wearer, preventing accurate recognition by the system. As an extension, we conduct physical experiments with masks.
@@
In our experiments, we create five different types of masks using various collection methods and fabrication processes. We categorize them from A to E based on their realism, with Mask A being directly collected through photography and created using 2D printing technology with paper material. Masks B and Masks C are created by scanning facial data using 3D scanning technology, with sandstone and paper pulp materials, respectively. Masks D and Masks E are also created using 3D scanning and silicone material, with Masks D and E having higher facial fit and elasticity.
@@
Subsequently, we test these masks on the Common Corruptions and Appearance Variations we design, in order to examine the characteristics of different masks and their impact in OOD scenarios. The results are shown in Tab. I.35, Tab. I.36, Tab. I.37, Tab. I.38, Tab. I.39 and Tab. I.40, Tab. I.41, Tab. I.42, Tab. I.43, Tab. I.44. The data reveals that face masks demonstrate material-dependent vulnerability patterns in OOD scenarios, exhibiting degradation trends distinct from real faces. This observation suggests a potential avenue for enhancing spoof detection.
@@ Appendix H More Visualization Results
Appendix I Statement on Data Sources and User Privacy
We would like to clarify that all images used in our experiments originate from publicly available open-source datasets that comply with the relevant data usage policies. No private or personally identifiable images were uploaded to any third-party commercial system or Vision-Language Models. Additionally, for the face images collected in our physical experiments, we obtained explicit consent from the participants for their data to be used in our research. Our study strictly adheres to ethical guidelines and privacy regulations to ensure the responsible handling of all data.
Figure I.1: Graded radar charts for each model on LFW-C.
Figure I.2: Graded radar charts for each model on LFW-V.
Figure I.3: Graded line charts for each corruption on LFW-C.
Figure I.4: Graded line charts for each corruption on LFW-V.
Figure I.5: RCE results on CFP-C.
Figure I.6: RVE results on CFP-V.
Figure I.7: Graded radar charts for each model on CFP-C.
Figure I.8: Graded radar charts for each model on CFP-V.
Figure I.9: Graded line charts for each corruption on CFP-C.
Figure I.10: Graded line charts for each corruption on CFP-V.
Figure I.11: RCE results on YTF-C.
Figure I.12: RVE results on YTF-V.
Figure I.13: Graded radar charts for each model on YTF-C.
Figure I.14: Graded radar charts for each model on YTF-V.
Figure I.15: Graded line charts for each corruption on YTF-C.
Figure I.16: Graded line charts for each corruption on YTF-V.
@@ Appendix I Statement on Data Sources and User Privacy
Table I.1: Accuracy of 19 FR models on LFW-C level 1.
@@
Table I.2: Accuracy of 19 FR models on LFW-C level 2.
@@
Table I.3: Accuracy of 19 FR models on LFW-C level 3.
@@
Table I.4: Accuracy of 19 FR models on LFW-C level 4.
@@
Table I.5: Accuracy of 19 FR models on LFW-C level 5.
@@
Table I.6: Accuracy of 19 FR models on LFW-V level 1.
@@
Table I.7: Accuracy of 19 FR models on LFW-V level 2.
@@
Table I.8: Accuracy of 19 FR models on LFW-V level 3.
@@
Table I.9: Accuracy of 19 FR models on LFW-V level 4.
@@
Table I.10: Accuracy of 19 FR models on LFW-V level 5.
@@
Table I.11: Accuracy of 19 FR models on CFP-C.
@@
Table I.12: Accuracy of 19 FR models on CFP-C level 1.
@@
Table I.13: Accuracy of 19 FR models on CFP-C level 2.
@@
Table I.14: Accuracy of 19 FR models on CFP-C level 3.
@@
Table I.15: Accuracy of 19 FR models on CFP-C level 4.
@@
Table I.16: Accuracy of 19 FR models on CFP-C level 5.
@@
Table I.17: Accuracy of 19 FR models on CFP-V.
@@
Table I.18: Accuracy of 19 FR models on CFP-V level 1.
@@
Table I.19: Accuracy of 19 FR models on CFP-V level 2.
@@
Table I.20: Accuracy of 19 FR models on CFP-V level 3.
@@
Table I.21: Accuracy of 19 FR models on CFP-V level 4.
@@
Table I.22: Accuracy of 19 FR models on CFP-V level 5.
@@
Table I.23: Accuracy of 19 FR models on YTF-C.
@@
Table I.24: Accuracy of 19 FR models on YTF-C level 1.
@@
Table I.25: Accuracy of 19 FR models on YTF-C level 2.
@@
Table I.26: Accuracy of 19 FR models on YTF-C level 3.
@@
Table I.27: Accuracy of 19 FR models on YTF-C level 4.
@@
Table I.28: Accuracy of 19 FR models on YTF-C level 5.
@@
Table I.29: Accuracy of 19 FR models on YTF-V.
@@
Table I.30: Accuracy of 19 FR models on YTF-V level1.
@@
Table I.31: Accuracy of 19 FR models on YTF-V level2.
@@
Table I.32: Accuracy of 19 FR models on YTF-V level3.
@@
Table I.33: Accuracy of 19 FR models on YTF-V level4.
@@
Table I.34: Accuracy of 19 FR models on YTF-V level5.
@@
Table I.35: Accuracy of 8 open-source models on Mask-A of corruptions.
@@
Table I.36: Accuracy of 8 open-source models on Mask-B of corruptions.
@@
Table I.37: Accuracy of 8 open-source models on Mask-C of corruptions.
@@
Table I.38: Accuracy of 8 open-source models on Mask-D of corruptions.
@@
Table I.39: Accuracy of 8 open-source models on Mask-E of corruptions.
@@
Table I.40: Accuracy of 8 open-source models on Mask-A of variations.
@@
Table I.41: Accuracy of 8 open-source models on Mask-B of variations.
@@
Table I.42: Accuracy of 8 open-source models on Mask-C of variations.
@@
Table I.43: Accuracy of 8 open-source models on Mask-D of variations.
@@
Table I.44: Accuracy of 8 open-source models on Mask-E of variations.